Joseph has 14 feet of rope. He wants to cut it into two pieces so that the
longer piece is one foot less than four times the length of the smaller
piece. What should be the length of the longer piece of rope?

Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52800)

.

Your setup equation is

    x + (4x-1) = 14.


where x is the shorter piece.


From the equation

    5x = 14 + 1 = 15.

     x          = 15/5 = 3.


ANSWER.  The shorter piece is 3 ft long.  The other piece is 14-3 = 11 ft long.

Solved.


Answer by Edwin McCravy(20059)

Joseph has 14 feet of rope. He wants to cut it into two pieces so that the
longer piece is one foot less than four times the length of the shorter
piece. What should be the length of the longer piece of rope?

In this piece of information:

the longer piece is one foot less than four times the length of the shorter piece

The longer piece is mentioned first and the shorter piece is mentioned second.
There are occasional exceptions, but the best rule to follow is: 

"Let x equal what is mentioned second."

So we let x = the length of the shorter piece.

the longer piece is one foot less than four times the length of the shorter piece

Length of longer piece = 4x - 1

matrix(5,1,Length, of, the, shorter, piece)),

The equation comes from



x + (4x - 1) = 14
  x + 4x - 1 = 14
      5x - 1 = 14
         + 1  + 1
     ------------
      5x     = 15
        5x/5 = 15/5
           x = 3 feet   <--length of shorter piece

Length of longer piece = 4x - 1 = 4(3) - 1 = 12 - 1 = 11 feet    


Edwin